OpenStudy (anonymous):

hi..need help Solve 1/3w=1/2

OpenStudy (anonymous):

w=3/2

OpenStudy (anonymous):

thank u can you show how you got it?? thanks

jhonyy9 (jhonyy9):

not is right 2/6w =3w/6w 2=3w w=2/3